a group of 100 people touring europe includes 42 who speak greek, 56 who speak german, and 15 who speak neither language. How many people in the group speak both Greek and German?
Found 2 solutions by addingup, ikleyn:
Answer by addingup(3677)

15 speak neither language: 100-15 = 85
42 Greek and 56 German: 42+56 = 98
98-85 = 13 speak both languages

Answer by ikleyn(52847)